# The Vital Role of Gradient Clipping in Byzantine-Resilient Distributed Learning

### v3

@@ Abstract
Byzantine-resilient distributed machine learning seeks to achieve robust learning performance in the presence of misbehaving or adversarial workers. While state-of-the-art (SOTA) robust distributed gradient descent (Robust-DGD) methods were proven theoretically optimal, their empirical success has often relied on pre-aggregation gradient clipping. However, the currently considered static clipping strategy exhibits mixed results: improving robustness against some attacks while being ineffective or detrimental against others. We address this gap by proposing a principled adaptive clipping strategy, termed Adaptive Robust Clipping (ARC). We show that ARC consistently enhances the empirical robustness of SOTA Robust-DGD methods, while preserving the theoretical robustness guarantees. Our analysis shows that ARC provably improves the asymptotic convergence guarantee of Robust-DGD in the case when the model is well-initialized. We validate this theoretical insight through an exhaustive set of experiments on benchmark image classification tasks. We observe that the improvement induced by ARC is more pronounced in highly heterogeneous and adversarial settings.
@@ 1 Introduction
Conventional distributed learning algorithms are known to be vulnerable to misbehaving workers that could behave unpredictably (Blanchard et al., 2017; Kairouz et al., 2021; Guerraoui et al., 2023). Misbehavior may result from software and hardware bugs, data poisoning, or malicious players controlling part of the network. In the parlance of distributed computing, misbehaving workers are referred to as Byzantine (Lamport et al., 1982). Due to the growing influence of distributed learning in critical public-domain applications such as healthcare (Nguyen et al., 2022) and finance (Long et al., 2020), the problem of robustness to misbehaving workers, a.k.a. robust distributed learning, has received significant attention (Blanchard et al., 2017; Yin et al., 2018; Farhadkhani et al., 2022; Karimireddy et al., 2022; Gorbunov et al., 2023; Allouah et al., 2023a; Farhadkhani et al., 2023).
Robust distributed learning algorithms primarily rely on robust aggregation, such as coordinate-wise trimmed mean (CWTM) (Yin et al., 2018), geometric median (GM) (Chen et al., 2017) and multi-Krum (MK) (Blanchard et al., 2017). Specifically, in robust distributed gradient descent (Robust-DGD), the server aggregates the workers’ local gradients using a robust aggregation method, instead of simply averaging them. This protects the learning procedure from erroneous gradients sent by misbehaving workers. Recent work has made significant improvements over these aggregation rules by incorporating a pre-aggregation step such as bucketing (Karimireddy et al., 2022; Gorbunov et al., 2023) and nearest-neighbor mixing (NNM) (Allouah et al., 2023a), to tackle gradient dissimilarity resulting from data heterogeneity. The learning guarantee of the resulting Robust-DGD has been proven to be optimal (Allouah et al., 2023b), i.e., it cannot be improved without additional assumptions under the standard heterogeneity model of $(G,B)$-gradient dissimilarity (Karimireddy et al., 2020).
Despite its theoretical tightness, the empirical success of Robust-DGD has unknowingly relied on pre-aggregation gradient clipping (Mhamdi et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023a). Specifically, clipping the gradients of the workers prior to aggregation has been observed to sometimes enhance the algorithm’s empirical performance in the presence of adversarial workers, as evidenced in Figure 1a. Yet, this improvement lacks a concrete explanation, raising the question of whether the observed benefits of clipping are merely anecdotal. This leads to the natural inquiry: Why, and when, does pre-aggregation clipping improve robustness?
(a) Sign-flipping (SF) (Allen-Zhu et al., 2020)
(b) Label-flipping (LF) (Allen-Zhu et al., 2020)
Figure 1: Impact of pre-aggregation clipping, specifically static clipping with $C=2$ and our adaptive clipping algorithm ARC, on Robust-DGD with aggregation scheme CWTM $\circ$ NNM, under the SF and LF adversarial attacks. We consider the MNIST (Deng, 2012) dataset distributed amongst $10$ honest (non-adversarial) workers with extreme heterogeneity, and there are three adversarial workers.
@@
ARC dynamically adjusts the clipping threshold as per the gradients sent by the workers and the fraction of adversarial workers to be tolerated. We demonstrate that integrating ARC into Robust-DGD consistently improves its empirical performance (see Figures 1 and 2a), while also preserving the convergence guarantee of the original Robust-DGD algorithm. Moreover, we show that when the model initialization is good, ARC provably improves the robustness of Robust-DGD. The benefits of ARC are more pronounced as the fraction of misbehaving workers approaches the system’s breakdown point [^1] Breakdown point refers the minimum fraction of adversarial workers that can break the system, making it impossible to guarantee a bound on the learning error Allouah et al. (2023b). and when the data across the workers is highly heterogeneous. Our key results are summarized below. Critical comparisons to prior work are deferred to Section 6.
(2) Improved empirical robustness. We conduct experiments on MNIST (Deng, 2012), Fashion-MNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2014), across various data heterogeneity settings and adversarial regimes. Our results demonstrate that ARC significantly enhances the performance of state-of-the-art Robust-DGD methods, particularly in scenarios with high data heterogeneity (Figure 2a) and a large number of adversarial workers (Figure 4b).
(3) Improved learning guarantee. We demonstrate that ARC possesses an additional property, we call Bounded Aggregation Output, that is not satisfied by classical robust aggregation methods. Specifically, ARC constrains the norm of an adversarial gradient by that of an honest (non-adversarial) gradient. Leveraging this property, we show that the lower bound established under data heterogeneity in Allouah et al. (2023b) can be circumvented using ARC, provided the honest workers’ gradients are bounded at model initialization. An empirical validation of this insight is shown in Figure 2b. Such model initialization is often satisfiable in practice (Glorot & Bengio, 2010), highlighting the practical relevance of ARC. When the model is arbitrarily initialized, ARC recovers the original convergence guarantee of Robust-DGD in the worst case (see Figure 2b).
@@ 2 Problem Statement and Relevant Background
A robust distributed learning algorithm aims to output a good model despite the presence of adversarial (a.k.a., Byzantine (Lamport et al., 1982)) workers in the system (Su & Vaidya, 2016; Yin et al., 2018; Allouah et al., 2023a). Specifically, the goal is to design a distributed learning algorithm that can tolerate up to $f$ adversarial workers, of a priori unknown identity, out of $n$ workers. Adversarial workers need not follow a prescribed algorithm, and can send arbitrary information to the server. In the context of distributed gradient descent, adversarial workers can send incorrect gradients to the server (Baruch et al., 2019; Allen-Zhu et al., 2020; Xie et al., 2020; Karimireddy et al., 2022). We let $\mathcal{H}\subseteq[n]$, with $\left\lvert{\mathcal{H}}\right\rvert=n-f$, denote the set of honest workers that do not deviate from the algorithm. The objective of the server is to minimize the average loss function of honest workers given by
$\mathcal{L}_{\mathcal{H}}(\theta)\coloneqq\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\sum_{i\in\mathcal{H}}\mathcal{L}_{i}(\theta),~{}\forall\theta\in\mathbb{R}^{d}.$
While we can find a minimum of $\mathcal{L}_{\mathcal{H}}(\theta)$ when $\mathcal{L}_{\mathcal{H}}$ is convex, in general however the loss function is non-convex, and the optimization problem is NP-hard (Boyd & Vandenberghe, 2004). Therefore, we aim to find a stationary point of $\mathcal{L}_{\mathcal{H}}$ instead, i.e., $\theta^{*}$ such that $\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}(\theta^{\ast})}\right\rVert=0$. Formally, we define robustness to adversarial workers by $(f,\epsilon)$-resilience (Allouah et al., 2023a).
@@
If a distributed learning algorithm $\mathcal{A}$ is $(f,\varepsilon)$-resilient then it can tolerate up to $f$ adversarial workers. It is standard in robust distributed learning to design robust algorithms using $f$ as a parameter (Su & Vaidya, 2016; Blanchard et al., 2017; Yin et al., 2018; Karimireddy et al., 2022; Gorbunov et al., 2023; Allouah et al., 2023a), $f$ being the maximum tolerable number of adversarial workers.
@@
Robust distributed gradient descent. To tolerate adversarial workers, we replace the averaging operation in the classic distributed gradient descent (DGD) method with a robust aggregation, e.g., coordinate-wise trimmed mean (CWTM) and median (CWMed) (Yin et al., 2018), geometric median (GM) (Chen et al., 2017), and multi-Krum (MK) (Blanchard et al., 2017). This yields Robust-DGD, presented in Algorithm 1, where robust aggregation protects the learning from incorrect gradients sent by the adversarial workers. The robustness of an aggregation method can be quantified by the following property of $(f,\kappa)$-robustness, proposed by Allouah et al. (2023a).
@@
$(f,\kappa)$-robustness [^2] Unifies $(\delta_{\text{max}},c)$-robustness (Karimireddy et al., 2021) and $(f,\lambda)$-resilience (Farhadkhani et al., 2022). encompasses several robust aggregations, including the ones mentioned above and more (e.g., see Allouah et al. (2023a)). It has been shown that an aggregation method is $(f,\kappa)$-robust only if $\kappa\geq\frac{f}{n-2f}$ (Allouah et al., 2023a). Importantly, the asymptotic error of Robust-DGD with $(f,\kappa)$-robust aggregation, where $\kappa\in\mathcal{O}(f/(n-2f))$, matches the lower bound (recalled in Lemma 2.3(Allouah et al., 2023b). Note that the aforementioned aggregation methods (CWTM, CWMed, GM and MK) attain optimal robustness when composed with the pre-aggregation scheme: nearest neighbor mixing (NNM) (Allouah et al., 2023a). [^3] CWTM has been shown to be optimal even without NNM. Specifically, we recall the following result.
@@
For $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, the composition $\mathbf{F}\circ\mathbf{NNM}$ is $(f,\kappa)$-robust with
$\kappa\leq\frac{8f}{n-f}\left(1+\left(1+\frac{f}{n-2f}\right)^{2}\right).$
@@
Thus, when $\kappa\in\mathcal{O}\left(f/(n-2f)\right)$ (and smaller than $1/B^{2}$), Robust-DGD is optimal, i.e., its error matches the lower bound recalled in Lemma 2.3, as soon as $T\geq\Delta_{o}/\gamma\kappa G^{2}$. [^4] For $f=0$, Lemma 2.3 recovers the convergence of DGD, provided that $\kappa$ is tight, i.e., $\kappa\in\mathcal{O}(f/(n-2f))$.
@@ 3 Adaptive Robust Clipping (ARC) and its Properties
Thus, if the clipping threshold is fixed, i.e., independent from the input vectors, pre-aggregation clipping does not preserve the robustness of the original aggregation. This fragility of such static clipping is also apparent in practice, as shown by our experimental study in Appendix F (and Figure 1).
Description of ARC. ARC is an adaptive clipping scheme that only makes use of the standard robustness parameter $\nicefrac{{f}}{{n}}$, i.e., the tolerable fraction of adversarial workers. ARC is adaptive in the sense that the clipping threshold $C$ is not fixed but depends on the input vectors. Specifically, ARC clips the largest $k=\left\lfloor{2(f/n)(n-f)}\right\rfloor$ vectors using a clipping parameter given by the norm of the $(k+1)$-th largest input vector. The overall scheme is formally presented in Algorithm 2, and its computational complexity is $\mathcal{O}(nd+n\log(n))$ (see Appendix A). Thus, pre-composing more computationally expensive schemes like NNM and multi-Krum, which have a complexity of $\mathcal{O}(dn^{2})$, with ARC does not introduce a significant overhead.
@@
Robustness Guarantee. We present below the preservation of robustness guaranteed by ARC. Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation rule and $\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})\coloneqq\mathbf{F}(\mathbf{ARC}(x_{1},\dots,x_{n}))$
@@
Proofs for the results presented in this section are deferred to Appendix B. Since $\kappa\geq\frac{f}{n-2f}$ (recalled in Section 2), Theorem 3.2 implies that $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,3\kappa\right)$-robust. In other words, ARC preserves the robustness of the original aggregation scheme. Therefore, a convergence result for Robust-DGD with ARC follows verbatim from Lemma 2.6, replacing $\kappa$ with $3\kappa$. Despite this multiplicative factor of 3, we observe in the next section that incorporating ARC consistently improves the empirical performance of classical aggregation methods. A more detailed theoretical explanation will be provided later in Section 5.1.
@@ 4 Empirical Evaluation
In this section, we delve into the practical performance of ARC when incorporated in Robust-DSGD (Algorithm 3 in Appendix D), an order-optimal stochastic variant of Robust-DGD (Allouah et al., 2023a). We conduct experiments on standard image classification tasks, covering different adversarial scenarios. We empirically test four aggregation methods when pre-composed with ARC. We also contrast these outcomes against the performance of Robust-DSGD when no gradient clipping is used. Our findings underscore that clipping workers’ gradients using ARC prior to the aggregation is crucial to ensure the robustness of existing algorithms, especially in extreme scenarios, i.e., when either the data heterogeneity is high or the fraction of adversarial workers is large.

### v5

# Adaptive Gradient Clipping for Robust Federated Learning

@@ Abstract
Robust federated learning aims to maintain reliable performance despite the presence of adversarial or misbehaving workers. While state-of-the-art (SOTA) robust distributed gradient descent (Robust-DGD) methods were proven theoretically optimal, their empirical success has often relied on pre-aggregation gradient clipping. However, existing static clipping strategies yield inconsistent results: enhancing robustness against some attacks while being ineffective or even detrimental against others. To address this limitation, we propose a principled adaptive clipping strategy, Adaptive Robust Clipping (ARC), which dynamically adjusts clipping thresholds based on the input gradients. We prove that ARC not only preserves the theoretical robustness guarantees of SOTA Robust-DGD methods but also provably improves asymptotic convergence when the model is well-initialized. Extensive experiments on benchmark image classification tasks confirm these theoretical insights, demonstrating that ARC significantly enhances robustness, particularly in highly heterogeneous and adversarial settings.
@@ 1 Introduction
Conventional distributed learning algorithms are known to be vulnerable to misbehaving workers that could behave unpredictably (Blanchard et al., 2017; Kairouz et al., 2021; Guerraoui et al., 2023). Misbehavior may result from software and hardware bugs, data poisoning, or malicious players controlling part of the network. In the parlance of distributed computing, misbehaving workers are referred to as Byzantine (Lamport et al., 1982). Due to the growing influence of distributed learning in critical public-domain applications such as healthcare (Nguyen et al., 2022) and finance (Long et al., 2020), the problem of robustness to misbehaving workers, a.k.a. robust distributed learning, has received significant attention (Yin et al., 2018; Farhadkhani et al., 2022; Karimireddy et al., 2022; Gorbunov et al., 2023; Allouah et al., 2023a; Farhadkhani et al., 2023; El-Mhamdi et al., 2023).
Robust distributed learning algorithms primarily rely on robust aggregation, such as coordinate-wise trimmed mean (CWTM) (Yin et al., 2018), geometric median (GM) (Chen et al., 2017) and Multi-Krum (MK) (Blanchard et al., 2017). Specifically, in robust distributed gradient descent (Robust-DGD), the server aggregates the workers’ local gradients using a robust aggregation method, instead of simply averaging them. This protects the learning from erroneous gradients sent by misbehaving workers. Recent work has made significant improvements over these aggregation techniques by incorporating a pre-aggregation step such as bucketing (Karimireddy et al., 2022; Gorbunov et al., 2023) and nearest-neighbor mixing (NNM) (Allouah et al., 2023a), to tackle gradient dissimilarity resulting from data heterogeneity. The learning guarantee of the resulting Robust-DGD has been proven optimal (Allouah et al., 2023b), i.e., it cannot be improved without additional assumptions under the standard heterogeneity model of $(G,B)$-gradient dissimilarity (Karimireddy et al., 2020).
Despite its theoretical tightness, the empirical success of Robust-DGD has unknowingly relied on pre-aggregation gradient clipping (Mhamdi et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023a). Specifically, clipping the gradients of the workers prior to aggregation has been observed to sometimes enhance the algorithm’s empirical performance in the presence of adversarial workers (as evidenced in Figure 7(a)). Yet, this improvement lacks a concrete explanation, raising the question of whether the observed benefits of clipping are merely anecdotal. This leads to the natural inquiry: Why, and when, does pre-aggregation clipping improve robustness?
@@
ARC dynamically adjusts the clipping threshold as per the gradients sent by the workers and the fraction of adversarial workers to be tolerated. We demonstrate that integrating ARC into Robust-DGD consistently improves its empirical performance (see Figures 7 and 1(a)), while also preserving the convergence guarantee of the original Robust-DGD algorithm. Moreover, we show that when the model initialization is good, ARC provably improves the robustness of Robust-DGD. The benefits of ARC are more pronounced as the fraction of misbehaving workers approaches the system’s breakdown point [^1] Breakdown point refers to the minimum fraction of adversarial workers that can break the system, making it impossible to guarantee a bound on the learning error (Allouah et al., 2023b). and when the data across the workers is highly heterogeneous. Our key results are summarized below. Critical comparisons to prior work are deferred to Section 6.
@@
(2) Improved empirical robustness. We conduct experiments on MNIST (Deng, 2012), Fashion-MNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2014), across various data heterogeneity settings and adversarial regimes. Our results demonstrate that ARC significantly enhances the performance of state-of-the-art Robust-DGD methods, particularly in scenarios with high data heterogeneity (Figure 1(a)) and a large number of adversarial workers (Figure 4(b)).
(3) Improved learning guarantee. We demonstrate that ARC possesses an additional property that is not satisfied by classical robust aggregation methods. Specifically, ARC constrains the norm of an adversarial gradient by that of an honest (non-adversarial) gradient. Leveraging this property, we show that ARC circumvents the lower bound established under data heterogeneity in Allouah et al. (2023b), provided the honest gradients are bounded at model initialization. An empirical validation of this insight is shown in Figure 1(b). Such model initialization is often satisfiable in practice (Glorot & Bengio, 2010), highlighting the practical relevance of ARC. When the model is arbitrarily initialized, ARC recovers the original convergence guarantee of Robust-DGD in the worst case.
@@ 2 Problem Statement and Relevant Background
A robust distributed learning algorithm aims to output a good model despite the presence of adversarial (a.k.a., Byzantine) workers in the system (Su & Vaidya, 2016; Yin et al., 2018; Allouah et al., 2023a). Specifically, the goal is to design a distributed learning algorithm that can tolerate up to $f$ adversarial workers, of a priori unknown identity, out of $n$ workers. Adversarial workers need not follow a prescribed algorithm, and can send arbitrary information to the server. In the context of distributed gradient descent, adversarial workers can send incorrect gradients to the server (Baruch et al., 2019; Xie et al., 2020). We let $\mathcal{H}\subseteq[n]$, with $\left\lvert{\mathcal{H}}\right\rvert=n-f$, denote the set of honest workers that do not deviate from the algorithm. The objective of the server is to minimize the average loss function of honest workers given by $\mathcal{L}_{\mathcal{H}}(\theta)\coloneqq\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\sum_{i\in\mathcal{H}}\mathcal{L}_{i}(\theta),~{}\forall\theta\in\mathbb{R}^{d}$. While we can find a minimum of $\mathcal{L}_{\mathcal{H}}(\theta)$ when $\mathcal{L}_{\mathcal{H}}$ is convex, in general however the loss function is non-convex, and the optimization problem is NP-hard (Boyd & Vandenberghe, 2004). Therefore, we aim to find a stationary point of $\mathcal{L}_{\mathcal{H}}$ instead, i.e., $\theta^{*}$ such that $\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}(\theta^{\ast})}\right\rVert=0$. Formally, we define robustness to adversarial workers by $(f,\epsilon)$-resilience (Allouah et al., 2023a).
@@
If a distributed learning algorithm $\mathcal{A}$ is $(f,\varepsilon)$-resilient, then it can tolerate up to $f$ adversarial workers. It is standard in robust distributed learning to design robust algorithms using $f$ as a parameter (Blanchard et al., 2017; Yin et al., 2018; Gorbunov et al., 2023; Allouah et al., 2023a).
@@
Robust distributed gradient descent. To tolerate adversarial workers, we replace the averaging operation in the classic distributed gradient descent (DGD) method with a robust aggregation, e.g., coordinate-wise trimmed mean (CWTM) and median (CWMed) (Yin et al., 2018), geometric median (GM) (Chen et al., 2017), and Multi-Krum (MK) (Blanchard et al., 2017). This yields Robust-DGD, presented in Algorithm 1, where robust aggregation protects the learning from incorrect gradients sent by the adversarial workers. The robustness of an aggregation method can be quantified by the following property of $(f,\kappa)$-robustness (Allouah et al., 2023a).
@@
$(f,\kappa)$-robustness encompasses several robust aggregations, including the ones mentioned above and more (e.g., see Allouah et al. (2023a)). It has been shown that an aggregation method is $(f,\kappa)$-robust only if $\kappa\geq\frac{f}{n-2f}$ (Allouah et al., 2023a). Importantly, the asymptotic error of Robust-DGD with an $(f,\kappa)$-robust aggregation, where $\kappa\in\mathcal{O}(f/(n-2f))$, matches the lower bound (recalled in Lemma 2.3(Allouah et al., 2023b). This testifies to the tightness of the $(f,\kappa)$-robustness property. Note that the aforementioned aggregation methods (CWTM, CWMed, GM and MK) attain optimal robustness when composed with the pre-aggregation scheme: nearest neighbor mixing (NNM) (Allouah et al., 2023a). Specifically, we recall the following result.
@@
For $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, the composition $\mathbf{F}\circ\mathbf{NNM}$ is $(f,\kappa)$-robust with $\kappa\leq\frac{8f}{n-f}\left(1+\left(1+\frac{f}{n-2f}\right)^{2}\right)$.
@@
Thus, when $\kappa\in\mathcal{O}\left(f/(n-2f)\right)$ (and smaller than $1/B^{2}$), Robust-DGD is optimal, i.e., its error matches the lower bound recalled in Lemma 2.3, as soon as $T\geq\Delta_{o}/\gamma\kappa G^{2}$. For $f=0$, Lemma 2.6 recovers the convergence of DGD, provided that $\kappa$ is tight, i.e., $\kappa\in\mathcal{O}(\frac{f}{n-2f})$.
@@ 3 Adaptive Robust Clipping (ARC) and its Properties
Thus, if the clipping threshold is fixed, i.e., independent from the input vectors, pre-aggregation clipping does not preserve the robustness of the original aggregation. This fragility of such static clipping is also apparent in practice, as shown by our experimental study in Section 4 and Appendix E.
Description of ARC. ARC is an adaptive clipping scheme that only makes use of the standard robustness parameter $f$, i.e., the tolerable number of adversarial workers. ARC is adaptive in the sense that the clipping threshold $C$ is not fixed but depends on the input vectors. Specifically, ARC clips the largest $k=\left\lfloor{2(f/n)(n-f)}\right\rfloor$ vectors using a clipping parameter given by the norm of the $(k+1)$-th largest input vector. The overall scheme is formally presented in Algorithm 2, and its computational complexity is $\mathcal{O}(nd+n\log(n))$ (see Appendix A.2). Thus, pre-composing more computationally expensive schemes like NNM and Multi-Krum, which have a complexity of $\mathcal{O}(dn^{2})$, with ARC does not introduce a significant overhead (see Appendix F.4). For a detailed explanation of the underlying intuition behind our adaptive clipping strategy, ARC, we refer the reader to Appendix A.1.
@@
Robustness Guarantee. We present below the preservation of robustness guaranteed by ARC. Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation method and $\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})\coloneqq\mathbf{F}(\mathbf{ARC}(x_{1},\dots,x_{n}))$
@@
Proofs for the results presented in this section are deferred to Appendix B. Since $\kappa\geq\frac{f}{n-2f}$ (recalled in Section 2), Theorem 3.2 implies that $\mathbf{F}\circ\mathbf{ARC}$ is $\left(f,3\kappa\right)$-robust, i.e., ARC preserves the robustness of the original aggregation scheme. Therefore, a convergence result for Robust-DGD with ARC follows verbatim from Lemma 2.6, replacing $\kappa$ with $3\kappa$. Despite this multiplicative factor of 3, we observe in Section 4 that incorporating ARC consistently improves the empirical performance of classical aggregation methods. A more detailed theoretical explanation is provided later in Section 5.1.
@@ 4 Empirical Evaluation
In this section, we delve into the practical performance of ARC when incorporated in Robust-DSGD (Algorithm 3 in Appendix D), an order-optimal stochastic variant of Robust-DGD (Allouah et al., 2023a). We conduct experiments on standard image classification tasks, covering different adversarial scenarios. We empirically test four aggregation methods when pre-composed with ARC. We contrast these outcomes against the performance of these aggregations under static clipping and when no gradient clipping is used. We show in Table 1, and Figures 2 and 4, the metric of worst-case maximal accuracy. For each of the five Byzantine attacks executed (see Appendix D.3), we record the maximal accuracy achieved by Robust-DSGD during the learning procedure under that attack. The worst-case maximal accuracy is thus the smallest maximal accuracy reached across the five attacks. Furthermore, we use the Dirichlet (Hsu et al., 2019a) distribution of parameter $\alpha$ to simulate data heterogeneity. The comprehensive experimental setup can be found in Appendix D. In this section, we focus on presenting our results for the CWTM and GM aggregation methods applied to MNIST and CIFAR-10. As benchmark, we also execute the standard DSGD algorithm, in the absence of adversarial workers (i.e., without attack and $f=0$). Results for Fashion-MNIST, as well as for CWMed and MK, are provided in Appendices E and  F. All aggregation methods in our experiments are pre-composed with NNM (Allouah et al., 2023a), but we omit explicit mention of NNM in their names for simplicity.
4.1 Brittleness of Static Clipping and Superiority of ARC
(a) $f=3$
(b) $f=4$
Figure 2: Impact of the clipping strategy, under varying heterogeneity levels, on the worst-case maximal accuracy achieved by Robust-DSGD on MNIST with $n=15$ workers. CWTM $\circ$ NNM is used as aggregation. DSGD reaches at least 98.5% in accuracy in all heterogeneity regimes.
We empirically demonstrate the brittleness of static clipping in robust distributed learning. While an exhaustive empirical study can be found in Appendix E, we present findings on MNIST using CWTM, considering three heterogeneity levels: moderate ($\alpha=1$), high ($\alpha=0.1$), and extreme. We execute Robust-DSGD with $n=15$ workers, among which $f\in\{3,4\}$ are adversarial, and test static clipping thresholds $C\in\{0.02,0.2,2,20\}$. First, we observe that the optimal $C$ depends heavily on data heterogeneity. As seen in Figure 2, $C=2$ is ideal in high heterogeneity, whereas $C=20$ performs much better under extreme heterogeneity. This dependency requires fine-tuning, which is impractical in distributed settings. Moreover, static clipping’s effectiveness varies significantly across Byzantine attacks. While $C=2$ performs well under SF (Allen-Zhu et al., 2020), it fails under LF, causing training collapse (see Figure 7 in Appendix E). This unpredictability makes static clipping unreliable in practice. The number of adversarial workers further affects static clipping’s performance. In Figure 2, $C=2$ is optimal for $f=3$ under high heterogeneity, yet its accuracy drops below 50% when $f=4$, making it ineffective. Unlike static clipping, ARC consistently matches or outperforms the best static threshold across all configurations, as shown in Figure 2. These results underscore the superiority of ARC over static clipping, eliminating the need for manual tuning and ensuring robustness regardless of unpredictable parameters like data heterogeneity and Byzantine attacks.
4.2 Performance Gains of ARC Over Robust-DSGD
To quantify the improvement induced by ARC over Robust-DSGD without clipping, we evaluate a distributed system with $n-f=10$ honest workers while varying the number of adversarial workers $f$. We compare the performance of Robust-DSGD with and without ARC on the MNIST dataset, highlighting the improvements achieved by our method. Additionally, we extend our analysis to larger distributed systems with 30 honest workers and $f\in\{3,6,9\}$, as detailed in Appendix F.1.2.
ARC boosts robustness in high heterogeneity. Figure 3 shows the performance of Robust-DSGD against the FOE attack when using ARC opposed to no clipping. In low heterogeneity (i.e., $\alpha=0.5$), the performances of ARC and no clipping are comparable. When the heterogeneity increases ($\alpha=0.1$), CWTM and GM significantly struggle to learn, while the same aggregations composed with ARC almost match the final accuracy of DSGD. In extreme heterogeneity, the improvement induced by ARC is the most pronounced, enabling both aggregations to reach a final accuracy close to 90%. Contrastingly, the same aggregations without clipping stagnate at around 10% throughout the training. Interestingly, only $f=1$ adversarial worker among $n=11$ (i.e., less than 10%) suffices to completely deteriorate the learning when no clipping is applied, highlighting the necessity of ARC in extreme heterogeneity. These observations are also conveyed in Figure 4(a) which compares the worst-case maximal accuracies achieved by Robust-DSGD when $f=1$, with and without ARC, for varying levels of heterogeneity. While CWTM and GM yield comparable accuracies with ARC in low heterogeneity ($\alpha\geq 0.5$), their performance significantly degrades when $\alpha$ drops below that threshold. Indeed, when $\alpha=0.1$, their accuracies drop below 65%, while ARC enables the same aggregations to maintain their accuracy at just below 98%. In extreme heterogeneity, the performance of the aggregations without clipping completely deteriorates with accuracies close to 15%. Contrastingly, ARC efficiently mitigates the Byzantine attacks, resulting in accuracies above 85% in the worst case for both aggregations. Similar plots for $f\in\{3,5,7\}$ convey similar observations (see Appendix F.1).